### v1

# FreshStack: Building Realistic Benchmarks for Evaluating Retrieval on Technical Documents

Abstract.
We introduce FreshStack, a reusable framework for automatically building information retrieval (IR) evaluation benchmarks from community-asked questions and answers. FreshStack conducts the following steps: (1) automatic corpus collection from code and technical documentation, (2) nugget generation from community-asked questions and answers, and (3) nugget-level support, retrieving documents using a fusion of retrieval techniques and hybrid architectures. We use FreshStack to build five datasets on fast-growing, recent, and niche topics to ensure the tasks are sufficiently challenging. On FreshStack, existing retrieval models, when applied out-of-the-box, significantly underperform oracle approaches on all five topics, denoting plenty of headroom to improve IR quality. In addition, we identify cases where rerankers do not clearly improve first-stage retrieval accuracy (two out of five topics). We hope that FreshStack will facilitate future work toward constructing realistic, scalable, and uncontaminated IR and RAG evaluation benchmarks. FreshStack datasets are available at: https://fresh-stack.github.io.
Information Retrieval, General Framework, Evaluation Benchmark
1. Introduction
Retrieval-augmented generation (RAG) is a popular technique to enhance traditional information retrieval (IR) capabilities with language model generation. RAG systems use large language models (LLMs) to generate long-form responses (Guu et al., 2020; Lewis et al., 2020; Izacard and Grave, 2021; Borgeaud et al., 2022), grounded in the information available from retrieved documents (Khandelwal et al., 2020; Lewis et al., 2020; Gao et al., 2023b; Liu et al., 2024a). Despite its wide usage, evaluating RAG remains incredibly challenging. Existing IR and RAG benchmarks are not well-suited for evaluation, as these are outdated and highly limited. In particular, we observe three major issues in existing benchmarks:
  • Lack of realistic, open-ended questions: Existing datasets contain purely extractive short answers (e.g., Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017)) or crowd-sourced questions (e.g., HotPotQA (Yang et al., 2018)). A limited number of datasets capture “natural” human-asked questions, i.e., MS MARCO (Nguyen et al., 2016) or Natural Questions (Kwiatkowski et al., 2019), but unfortunately, brief and straightforward questions are inserted into a search box, failing to represent the complex questions that real users might pose to modern RAG systems.
  • Artificially easy: RAG represents an approach rather than a problem. Real users require systems capable of grounded question answering, i.e., responding to specialized questions by referencing knowledge from a document corpus. Consequently, datasets constructed by design to be solvable via retrieval often fail to encode challenges faced in RAG applications.
  • Static and unspecialized: After sourcing questions and answers, a benchmark becomes at the risk of (1) contamination, if current LLMs are trained on the same set of documents or questions, (2) overfitting, when systems inevitably saturate by repeated internal or external leaderboarding (e.g., BEIR (Thakur et al., 2021)), and (3) staleness, when questions or answers are not refreshed and become outdated.
Figure 1. A data instance from LangChain generated with FreshStack. The question and answer pair is sourced from Stack Overflow and provided as input to generate nuggets with GPT-4o (highlighted in blue). Next, code snippets and technical documents from multiple GitHub repositories (e.g., Jupyter Notebook) are chunked, processed, and pooled for each question. Finally, each pooled document chunk is judged with GPT-4o for binary relevance (either yes or no) at a nugget-level, i.e., whether the document factually supports the information present in each nugget.
IR / RAG Benchmarks Niche Complex Dynamic Challenge
Domains Questions Updates Level
CQADupstack (Hoogeveen et al., 2015) No No No Easy
CodeSearchNet (Husain et al., 2019) No No No Easy
COIR (Li et al., 2024a) Limited Yes No Moderate
Stack Overflow-QA (Li et al., 2024a) No Yes No Moderate
CodeRAG-Bench (Wang et al., 2024a) Limited No No Moderate
Neural Code Search (Li et al., 2019) No No No Moderate
SWE-Bench (Jimenez et al., 2024) No Yes Yes High
FreshStack (ours) Yes Yes Yes High
Table 1. A comparison of existing IR/RAG evaluation benchmarks with FreshStack.
A realistic benchmark should measure model generalization on niche domains, and continue to update. Additionally, it must capture the complexity of human-generated queries—such as multi-hop reasoning (Yang et al., 2018), code understanding (Jimenez et al., 2024), or specialized terminology—rather than relying on artificially easy questions. This drives the robustness of systems in answering questions in evolving public libraries or private code bases (Zhang et al., 2023a; Jimenez et al., 2024), a company’s internal forum (Raina and Gales, 2024), or technical troubleshooting (Soman and Roychowdhury, 2024; Pu et al., 2025; Chen et al., 2025).
In our work, we introduce FreshStack, a holistic framework for constructing realistic datasets on niche and challenging domains, seeking to avoid contamination due to (perpetual) recency. Using FreshStack, we construct an evaluation benchmark on five niche topics sourced from community-asked questions and answers on Stack Overflow and a corpus containing code snippets and technical documents from public GitHub repositories. The FreshStack framework contains three major steps:
  1. (1)
    Automatic corpus collection (Section 3.2) with technical documents chunked and sourced from several GitHub repositories.
  2. (2)
    Nugget generation (Section 3.3) with GPT-4o using community-asked questions and answers in Stack Overflow.
  3. (3)
    Nugget-level support (Section 3.5) with GPT-4o on document chunks, retrieved from a fusion of retrieval techniques and hybrid architectures.
FreshStack has several advantages over existing benchmarks (as highlighted in Table 1). First, the framework utilizes user-asked questions and curated answers, making the evaluation challenging. We are not crafting artificial (LLM-generated) questions or sampling questions myopically. Second, all answers are supported in real time by information from technical documentation in GitHub repositories. Third, the framework is designed to be general and scalable without modification. Finally, the framework is focused on niche domains and recent topics, taking careful measures to mitigate risks with data contamination introduced by LLMs.
We provide a complete overview of our framework and investigate three research questions (RQs) to provide insights into the design choices used in FreshStack:
  • RQ1
    How to construct realistic and challenging IR evaluation datasets from user-asked questions?
  • RQ2
    How do LLMs act as an assessor for (1) nugget generation with community-asked questions & answers and (2) nugget-level support with GitHub-sourced documents?
  • RQ3
    How do out-of-the-box retrieval and reranker systems perform on benchmark generated with FreshStack?
We investigate the framework’s quality by constructing five datasets on niche topics, e.g., LangChain or Godot. We calibrate the automatic stage with GPT-4o using a machine learning (ML) expert, assessing the quality of nugget generation and nugget-level support for one of the topics (LangChain). Our results show that LLM-generated nuggets capture crucial information required to answer the question, and GPT-4o precisely labels support at a nugget level.
For the judgment pool construction, we assess pooling results by comparing the oracle (having access to the answer) and inference (relying only on the question) settings, showing that ensemble fusion in question decomposition (in the inference setting) and nugget generation (in the oracle setting) outperforms other techniques, respectively. Beyond pool construction, we explore the document retrieval setting on FreshStack, evaluating retrieval and rerankers out-of-the-box, using the question alone. We find that retrieval models drastically underperform oracle systems on all five topics, showing a high headroom for improvement on specialized topics. In addition, ensemble fusion outperforms individual models, indicating that diversity in models enhances retrieval, and rerankers provide clear benefits in some but not all topics.
FreshStack is a general framework and can be applied to any domain of a similar structure. The datasets will be made publicly available to the research community. Overall, we hope FreshStack can serve as a testbed for future work to develop similar challenging benchmarks for evaluating IR and RAG systems.
2. Related Work
Retrieval-augmented generation. RAG has been widely used to avoid “hallucinations” (Zhang et al., 2023b) seen by LLMs when handling knowledge-intensive tasks (Kandpal et al., 2023). RAG reduces the generation of factually incorrect content, leading to adoption in various commercial systems, e.g., Bing Search or Google AI Overviews. Existing IR and RAG benchmarks are stale, focused on text, evaluating on academic question answering datasets (Gao et al., 2023b; Rosenthal et al., 2025; Ram et al., 2023), or are not challenging, being constructed for RAG (Yang et al., 2024; Chen et al., 2024; Li et al., 2024b; Niu et al., 2024). A limited number of datasets refresh over time to avoid LLM decontamination (Kasai et al., 2023; Vu et al., 2024; Shao et al., 2024; White et al., 2025), however, these contain easy and unrealistic questions for evaluation. In contrast, FreshStack generates niche and challenging datasets, which are not constructed for RAG, and can refresh over time.
Code-based benchmarks. Neural code generation (Lu et al., 2021) requires LLMs to generate code from scratch for generic programming questions. One popular benchmark is SWE-Bench (Jimenez et al., 2024), which evaluates whether LLMs can generate code changes for GitHub pull requests (PRs) in popular public repositories. Similarly, CodeSearchNet (Husain et al., 2019), COIR (Li et al., 2024a), LiveCodeBench (Jain et al., 2024), and CodeRAG-Bench (Wang et al., 2024a) focus on the evaluation of high-level programming problems on popular public repositories. In contrast, in FreshStack, we focus on assisting developers, from a novice to a domain expert, by providing real-time answers on fast-growing and recent topics such as LangChain (introduced in 2023) by referencing technical documentation in GitHub repositories.
Stack Overflow datasets. FreshStack is not the first dataset to use Stack Overflow for retrieval. However, the evaluation setting of retrieving canonical documents from GitHub repositories remains under-explored. Existing datasets such as CQADupstack (Hoogeveen et al., 2015), LoTTE (Santhanam et al., 2022), and Stack Overflow-QA (Li et al., 2024a) follow a different task, to retrieve the relevant answer snippet given the question asked by a real user on Stack Overflow. The closest setting similar to FreshStack is found in Neural Code Search (Li et al., 2019), which incorporates public documentation and code snippet examples from GitHub as the corpus to answer questions asked by real users on popular programming topics such as Android.
Figure 2. The FreshStack framework: (1) Stack Overflow questions and answers are sourced for recent and niche topics. (2) GitHub repository documents are collected and chunked to form the corpus (for each topic). (3) Nuggets or atomic facts within the question and answer are generated with GPT-4o. (4) Ensemble techniques and models retrieve documents, which construct our document judgment pools. (5) GPT-4o evaluates support for every document-nugget pair as a binary judgment.
3. The FreshStack Framework
The FreshStack framework involves five stages to construct an end-to-end evaluation dataset (highlighted in Figure 2). The framework includes three major design choices:
  1. (1)
    Inclusion of recent and niche topics actively discussed in computer programmer communities such as Stack Overflow.
  2. (2)
    General and automatic framework that can be extended to different domains (potentially even across languages) without much manual effort.
  3. (3)
    Sourcing community-asked questions and answers, to make our evaluation challenging, requiring domain expert knowledge to answer them correctly.
Stack Overflow is an online question answering platform for computer programmers. It functions as a collaborative knowledge base, utilizing a voting system to curate content quality. Users ask questions about a particular topic and provide a description (often a code snippet with the error message) with a Stack Overflow tag. Anyone in the community can answer the question, however, higher-voted answers are prioritized, reflecting community consensus on the accuracy and relevance of the answer. Questions and answers are also tagged by topics, allowing for easy retrieval of topic-wise questions.
3.1. Stack Overflow Topic and Question Selection
For topic selection, we target niche and recent topics introduced on Stack Overflow from 2023 onward, containing a minimum of 50 posts. We sort all topics using the overall number of posts and curate five topics starting from the highest (LangChain) to the lowest frequency (Yolo v7 & v8) covering different domains and sufficiently different from each other.
Questions & Answers. We extract relevant posts and answers from the Stack Overflow XML data archive (dated October 2024). [^1] The Stack Overflow XML data archive (CC BY-SA license) is updated once every quarter: https://meta.stackexchange.com/questions/401324/announcing-a-change-to-the-data-dump-process We scan the archive to pick all the relevant posts as questions containing the required tag and to filter answer posts to the questions. At the end, we filter and keep questions with an accepted answer, prioritizing precision over quantity in retaining questions with high-quality answers.
3.2. Automatic Corpus Collection with GitHub
Answering a higher percentage of questions requires a robust set of corpora from multiple sources. For instance, addressing issues in LangChain may require ChromaDB GitHub documentation to resolve errors related to its usage. In our work, we build a different corpus set per topic by combining multiple GitHub repositories as sources (we list the GitHub repositories per topic in Table 2).
Table 2. GitHub repositories used for constructing the document collection for every topic in FreshStack.
Stack Overflow Tags. We analyze which GitHub repository to select for the document corpus by analyzing tag frequency from Stack Overflow posts. This involves identifying top-k co-occurring tags, where k is the threshold balancing question coverage with indexing costs. Some tags are generic, such as Python, whereas others are specific, such as LangChainJS. Filtering the tags to keep only a subset of repositories does not degrade the FreshStack dataset quality. We manually verify GitHub repositories for each tag, with plans to automate this procedure in the future. We could have constructed the corpus directly, however, the top-k co-occurring tags provide a useful signal on relevant GitHub repositories.
Chunking & Indexing. We clone the latest branch of a GitHub repository in a local workspace and parse all files as a tree structure. Each file (either a text document or code snippet) is chunked into small sections of up to a maximum of 4096 tokens, skipping non-text formats. [^2] We skip indexing images, videos, .bin, .csv, and audio files or unrecognized file formats. The GitHub filepath serves as the document identifier, with additional chunk details encoded in the identifier. Finally, we combine all chunks into a single corpus, prefixing all document identifiers with the repository name to identify the common files in each repository separately (e.g., LICENSE or requirements.txt).
3.3. Nuggetization or Nugget Generation
A nugget is a core concept or atomic fact essential in a system’s response. The term nugget was informally referred to as SCU (summary content units) as clauses appearing in model summarization (Nenkova and Passonneau, 2004) and later formalized as “information nugget” for evaluating long-form answers (Voorhees, 2003; Lin and Demner-Fushman, 2005, 2006; Pavlu et al., 2012). Nuggetization refers to constructing or generating nuggets from information-dense text. The procedure decomposes a verbose answer into key atomic facts or essential components, aiding evaluation. More recently, with the onset of RAG, nugget-based evaluation has renewed interest with LLMs for factual accuracy assessment in long answers (Arabzadeh and Clarke, 2024; Mayfield et al., 2024; Farzi and Dietz, 2024; Pradeep et al., 2024).
Nuggetization. We automatically generate nuggets from Stack Overflow question-answer pairs using GPT-4o (OpenAI, 2024), avoiding the cumbersome procedure of manual nugget construction (Pradeep et al., 2024). LLM-based nugget generation is explored in the TREC 2024 RAG track [^3] TREC 2024 RAG track: https://trec-rag.github.io/ (Pradeep et al., 2024) and in multiple works (Dietz, 2024; Farzi and Dietz, 2024). Separately, we experimented with prompting techniques and found that grading notes style prompts (Liu et al., 2024b) provided parseable and high-quality nuggets in our experiments.
3.4. Retrieval: Oracle & Inference Setting
A RAG evaluation dataset requires questions, answers, and a corpus with documents, which helps support facts in the answer. In this stage, we retrieve a list of highly relevant unjudged documents from the corpus and construct judgment pools. Since, we are constructing an evaluation dataset and we have curated answers for questions, we retrieve a list of top-k documents using two methods: (1) Inference, relying only on the question and automatic approaches, and (2) Oracle, relying on the gold answer or list of nuggets, to pool diverse documents for relevance judgment in the next stage.
Retrieval Settings. We experiment with multiple settings to increase diversity in our judgment pools. First, we experiment with two techniques in the inference setting:
  • GPT-4o Sub-Questions. Decomposing the original question, we generate a few synthetic sub-questions using GPT-4o, similar to Rosset et al. (2024), concatenated together to retrieve documents.
  • GPT-4o Closed Book Answer. We generate a closed-book answer for the original question with GPT-4o, similar to HyDE (Gao et al., 2023a), and use the closed-book answer to retrieve documents.
Next, in the Oracle setting, we also experiment with two techniques:
  • Stack Overflow Answer: We use the curated Stack Overflow answer as the question to retrieve documents.
  • Stack Overflow Nuggets: We use the list of GPT-4o generated nuggets (Section 3.3), concatenated as the question to retrieve documents.
Retrieval Models. We experiment with five different code and text-aware retrieval models:
  1. (1)
    BM25, a strong lexical baseline in BEIR (Thakur et al., 2021). We utilized the default BM25 implementation available in Pyserini (Lin et al., 2021).
  2. (2)
    BGE (Gemma-2) (Chen et al., 2024) a dense retriever model [^4] BGE Gemma-2: https://huggingface.co/BAAI/bge-multilingual-gemma2 fine-tuned on the backbone architecture of Gemma-2 (9B) (Rivière et al., 2024) with an embedding size of 3584 and 8K context length.
  3. (3)
    E5 Mistral (7B) (Wang et al., 2024b) is a dense retriever model [^5] E5 Mistral 7B: https://huggingface.co/intfloat/e5-mistral-7b-instruct based fine-tuned on the backbone of Mistral 7B (Jiang et al., 2023) with 32 layers and embedding size of 4096.
  4. (4)
    Voyage-large-2666Voyage-large-2: https://docs.voyageai.com/docs/embeddings is a proprietary and general-purpose embedding model optimized for retrieval quality, with a context length of 16K tokens and embedding size of 1536.
  5. (5)
    Fusion, a hybrid retrieval strategy with the four individual models, by normalizing and summing up the top 100 documents and their scores from each model.
3.5. Nugget-Level Support Assessment
This is the final stage in the framework. Traditionally, relevance judgments are conducted on selected pools of retrieved documents, i.e., a human assessor judges the relevance of the question and each provided document pair. Due to computational costs, recent studies experiment with an LLM-as-a-judge (instead of a human assessor) for conducting relevance judgments in information retrieval (Faggioli et al., 2023; Thomas et al., 2024; Upadhyay et al., 2024b, a; Rahmani et al., 2024). Questions in existing IR datasets are traditionally short, making it easier to assess question-document relevance. In contrast, questions in the FreshStack dataset are long and elaborate (between 350–500 tokens in length), containing a mixture of text, code snippets, or outputs, making it challenging to judge question-document relevance directly (Damessie et al., 2016). For instance, retrieved documents may answer a major problem presented in the question, address only part of the question, or contain relevant references and examples, and we need to translate this into a relevance score.
Topic Domain Dataset Count Avg. Length % Containing Code Relevance Judgments
#Queries #Docs #GitHub Avg. N/Q Query Answer Query Answer Rel. Docs/N Rel. Docs/Q
LangChain Machine Learning (ML) 203 49,514 10 3.1 473.4 233.8 83.3% 62.1% 5.7 10.9
Yolo v7 & v8 Computer Vision (CV) 57 27,207 5 3.5 497.1 191.7 70.2% 71.9% 3.9 7.4
Laravel 10 & 11 Backend Development 184 52,351 9 3.0 474.4 155.5 43.5% 51.1% 3.2 6.0
Angular 16, 17 & 18 Front-end Development 129 117,288 4 3.2 463.3 215.1 69.8% 57.4% 4.4 8.7
Godot4 Game Development 99 25,482 6 3.3 350.4 263.0 52.5% 52.5% 2.9 5.9
Table 3. FreshStack dataset statistics; Dataset count measures the number of queries, documents, GitHub repositories, and average nuggets per query; Avg. length measures the average text lengths (in tokens); % containing code measures the fraction of queries and answers with code snippets; Relevance judgments measure relevant documents per nugget and query.
Nugget-level Support. Instead of relying on traditional relevance assessments, we simplify the judgment procedure for the LLM and evaluate whether a document supports information (or contains) provided by a nugget. A reminder that a nugget highlights an essential fact of the Stack Overflow question or answer. Judging document relevance at a nugget level is effective as nuggets are factual and short information snippets, reducing the ambiguity often seen during traditional relevance judgments. As we have $n$ nuggets per question, repeating the assessment of $k$ documents for every nugget is prohibitive ($n\times k$). To reduce computational costs, we evaluate top-k documents (a maximum $k=20$) together with the list of all nuggets for a question in a single inference call ($n+k$). We evaluate support judgment with GPT-4o using chain-of-thought prompting (Wei et al., 2022).
4. Dataset Statistics & Evaluation
Upon completion of previous stages, we employ two post-processing steps to ensure high-quality question and answer pairs remain in the dataset, reducing the overall dataset size. In the first post-processing stage, we remove unsupported questions, i.e., questions that do not contain even a single relevant document; this removes, on average, 11.8% of the total questions. [^7] Future work may include these questions as they are potentially valuable to answer, and better retrieval systems may be able to find relevant documents. In the next stage, we aggressively filter by removing questions containing at least one unsupported nugget, i.e., a nugget that is not supported by any documents, reducing on average 34.2% of the total questions.
4.1. Dataset Statistics
FreshStack datasets covers five commonly used topics for programmers: machine learning, computer vision, backend, front-end, and game development, all listed in Table 3. Stack Overflow topics such as LangChain were introduced in 2023, whereas others, like Laravel or Angular, have questions about the latest versions (e.g., Laravel 10, Laravel 11). Each topic has at least 50 questions, and the corpus has at least 25K documents sourced from 4–10 GitHub repositories. The questions are long, containing 350–500 tokens (computed using GPT-4o tokenizer), and at least 50% of the questions and answers contain code snippets. GPT-4o generates around 3–4 nuggets for each question. Each nugget is matched to at least 3 relevant documents with support judgment from GPT-4o, resulting in 5–6 relevant documents per question, across all topics.
4.2. Instance Description
Each FreshStack dataset instance contains the following four components, as shown in Figure 1:
  • Question & Answer: The title and body (description) of the Stack Overflow post as the question, with the accepted answer. The title is a short sentence, and the body contains the detailed issue with code snippets and/or outputs.
  • Nuggets: The list of atomic facts highlighting the essential information in the Stack Overflow question and answer.
  • Document Corpus: The exhaustive list of chunked source documents (code snippets, text documentation, etc.) compiled from GitHub repositories.
  • Relevance Judgments: Unlike traditional IR benchmarks, such as BEIR (Thakur et al., 2021), which contain question and document-level relevance judgments, FreshStack datasets contain nugget-level relevance judgments for document chunks.
Method Model LangChain (ML) Yolo v7 & v8 (Vision) Laravel 10 & 11 (Backend) Angular 16, 17 & 18 (Front-end) Godot4 (Game Dev)
$\alpha$-N@10 Cov@20 Rec@50 $\alpha$-N@10 Cov@20 Rec@50 $\alpha$-N@10 Cov@20 Rec@50 $\alpha$-N@10 Cov@20 Rec@50 $\alpha$-N@10 Cov@20 Rec@50
Inference Setting: Using a variant of the Stack Overflow question for retrieval of documents within the corpus
GPT-4o Sub - Questions BM25 0.228 0.495 0.249 0.150 0.427 0.328 0.349 0.656 0.464 0.307 0.666 0.378 0.154 0.326 0.211
BGE (Gemma-2) 0.220 0.561 0.324 0.220 0.554 0.367 0.407 0.727 0.585 0.360 0.707 0.459 0.240 0.532 0.382
E5 Mistral (7B) 0.262 0.613 0.362 0.266 0.593 0.484 0.306 0.643 0.528 0.305 0.617 0.397 0.220 0.461 0.349
Voyage-large-2 0.270 0.563 0.329 0.213 0.526 0.370 0.366 0.687 0.552 0.344 0.69 0.449 0.260 0.594 0.473
Fusion (4 models) 0.322 0.708 0.475 0.305 0.665 0.489 0.478 0.763 0.662 0.428 0.817 0.584 0.290 0.598 0.526
GPT-4o Closed Book Answer BM25 0.256 0.520 0.273 0.286 0.554 0.431 0.376 0.655 0.495 0.293 0.542 0.332 0.241 0.473 0.349
BGE (Gemma-2) 0.181 0.467 0.263 0.271 0.599 0.473 0.36 0.694 0.539 0.242 0.525 0.338 0.187 0.454 0.358
E5 Mistral (7B) 0.198 0.471 0.277 0.239 0.511 0.364 0.188 0.458 0.384 0.179 0.430 0.267 0.151 0.318 0.237
Voyage-large-2 0.220 0.500 0.301 0.247 0.557 0.495 0.317 0.658 0.524 0.227 0.461 0.338 0.253 0.510 0.454
Fusion (4 models) 0.275 0.630 0.432 0.356 0.686 0.578 0.420 0.738 0.641 0.290 0.582 0.470 0.288 0.538 0.492
Oracle Setting: Using the Stack Overflow answer directly or its variants for retrieval of documents within the corpus
Stack Overflow Answer BM25 0.461 0.726 0.428 0.481 0.756 0.574 0.511 0.774 0.588 0.469 0.751 0.521 0.325 0.565 0.397
BGE (Gemma-2) 0.290 0.625 0.367 0.390 0.815 0.604 0.472 0.814 0.675 0.346 0.690 0.481 0.341 0.718 0.561
E5 Mistral (7B) 0.331 0.671 0.430 0.315 0.683 0.509 0.26 0.634 0.488 0.291 0.570 0.412 0.277 0.546 0.434
Voyage-large-2 0.385 0.700 0.432 0.405 0.703 0.589 0.439 0.791 0.641 0.371 0.680 0.477 0.371 0.626 0.541
Fusion (4 models) 0.484 0.821 0.619 0.546 0.854 0.788 0.564 0.892 0.820 0.470 0.805 0.695 0.449 0.741 0.683
Stack Overflow Nuggets BM25 0.467 0.739 0.445 0.519 0.796 0.657 0.540 0.840 0.654 0.485 0.787 0.536 0.428 0.680 0.489
BGE (Gemma-2) 0.308 0.667 0.405 0.461 0.784 0.572 0.448 0.806 0.666 0.393 0.756 0.536 0.335 0.664 0.555
E5 Mistral (7B) 0.323 0.684 0.432 0.437 0.737 0.554 0.287 0.631 0.533 0.346 0.670 0.470 0.292 0.596 0.494
Voyage-large-2 0.419 0.763 0.508 0.430 0.845 0.675 0.409 0.791 0.624 0.406 0.733 0.533 0.353 0.715 0.590
Fusion (4 models) 0.519 0.881 0.655 0.601 0.876 0.825 0.566 0.888 0.818 0.544 0.881 0.756 0.476 0.814 0.719
Table 4. Pooling results achieved by retrieval baselines (including fusion) in inference or oracle settings for constructing the FreshStack dataset. $\alpha$-N@10 denotes $\alpha$-nDCG@10, Cov@20 denotes Coverage@20 and Rec@50 denotes Recall@50. Stack Overflow Answer & Answer Nuggets methods rely on the gold answer for retrieval (oracle setting), whereas, other methods do not rely on the gold answer for retrieval. Overall, we highlight the best result in bold for each setting.
4.3. Retrieval Evaluation Metrics
Information retrieval (IR) evaluation follows the Cranfield paradigm (Voorhees, 2009), focusing on individual document relevance, independent of other documents. This is used to construct standard test collections, such as BEIR (Thakur et al., 2021) and TREC datasets such as the Deep Learning (DL) track (Craswell et al., 2022, 2023, 2024). However, diversity in search (Carbonell and Goldstein, 1998; Santos et al., 2015; Wu et al., 2024) penalizes information redundancy within retrieved documents to enrich information content and improve efficiency. Therefore, we evaluate retrieval systems using three diversity-focused metrics: $\alpha$-nDCG@10 for diversity and relevance, Coverage@20 for nugget coverage, and Recall@50 for traditional relevancy.
$\alpha$-nDCG@k. Introduced by Clarke et al. (2008), this variant of Normalized Discounted Cumulative Gain (nDCG) measures search diversification. The $\alpha$ parameter is a geometric penalization for redundant documents, i.e., each redundant document achieves a penalization of ×$\times(1-\alpha)$. Despite the metric being used for different user intents, we utilize it to ensure document rankings reference diverse nuggets in the answer.
Coverage@k. The metric introduced in our work measures the average proportion of the nuggets covered by the top-k retrieved documents. The mathematical formula is calculated as:
[…]
where $Q$ contains all questions, $\text{Nuggets}(d_{qi})$ are nuggets supported by document $d_{qi}$ and $\text{Nuggets}(q)$ are nuggets for question $q$.
Recall@k. The standard relevance metric measures the proportion of relevant documents retrieved within the top-k results, out of all relevant documents for a given question. A document is judged relevant if it supports at least one nugget.
5. Pooling & Qualitative Evaluation
Using FreshStack, we construct five IR datasets focused on document retrieval evaluation with a subset of question-answer pairs. [^8] For each question in the FreshStack dataset, the curated answer is present, allowing us to extend answer evaluation in RAG as potential future work. In this section, we attempt to answer RQ2 by evaluating techniques that contribute to judgment pools. Next, we calibrate the quality of GPT-4o against an expert human annotator. We first evaluate the retrieval baselines during nugget-level support judgment (or sampling pools) with inference and oracle settings.
In FreshStack, we are constructing a test evaluation dataset. Therefore, we can use the Stack Overflow answer or its variants in constructing judgment pools, as discussed previously in Section 3.4. We pool and sample documents from different settings and techniques, similar to how existing question answering datasets are constructed, such as Natural Questions (Kwiatkowski et al., 2019) or XOR-TyDI (Asai et al., 2021), which assess relevance by calculating the answer overlap in the document.
Experimental Settings. We perform retrieval with four techniques and baselines (explained in Section 3.4) and an ensemble fusion of the top 100 documents, with each model score normalized and summed up. Evaluation metrics include $\alpha$-nDCG@10, Coverage@20, and Recall@50. We use GPT-4o with a temperature setting of 0.1999Separately, we tested temperatures of 0.1 and 0.7, observing an identical downstream retrieval accuracy during FreshStack construction. for both the automatic stages. Nugget generation uses a grading notes prompt with the question and answer, and support assessment uses chain-of-thought prompting (Wei et al., 2022), judging up to a maximum of 20 documents simultaneously with a list of nuggets generated for each question. Finally, we sample and judge the top 20 fusion documents from each technique and setting, including the question to avoid sampling holes, highlighting the importance of document diversity in our judgment pools.
5.1. Pooling Results
We outline the results achieved on document judgment pools during FreshStack dataset construction with four different sampling techniques and models, including inference and oracle settings. Table 4 shows results for all five topics. Key takeaways and findings are discussed below:
Overall Highlights. Table 4 reveals two key findings: (1) Techniques in the Oracle setting significantly outperform techniques from the inference setting. We observe that both the Stack Overflow answer and nuggets techniques help pool documents relevant to the question, and (2) Fusion outperforms all individual models, highlighting the value of diversity in model choice, aiding in our judgment pools.
Inference Setting. GPT-4o Sub-Questions achieves the best pooling results for four topics (except Yolo v7 & v8), showing that decomposing a Stack Overflow question into smaller sub-questions helps pool documents relevant to the question. GPT-4o closed-book answer excels only on Yolo v7 & v8, which we suspect due to possible data contamination. These overall confirm that FreshStack contains non-contaminated data, as GPT-4o closed-book answer (For example, the ChatGPT generated answer) cannot retrieve relevant documents for answering niche questions in the FreshStack dataset on most topics.
Oracle Setting. Stack Overflow Nuggets achieves the best results on all topics (except Laravel 10 & 11), showing that GPT-4o-generated nuggets help pool more documents relevant to the question, instead of using the Stack Overflow answer. Amongst the individual models, BM25 achieves the best $\alpha$-nDCG@10 on all topics in both oracle methods (except Godot4), which asserts the importance of lexical approaches required during the construction of judgment pools in FreshStack.
5.2. Qualitative Analysis
In our work, a crucial component is the automatic construction of nuggets and nugget-level document judgments with GPT-4o. To assess the LLM quality, we calibrate using an expert annotator (ML researcher) conducting a qualitative analysis on a small subset of LangChain, evaluating the quality of generated nuggets and nugget-document support labels for 60 randomly sampled questions.
Nugget Quality Judgment Quality
Precision 90.1 % Relevant 71.7 %
Recall 96.6 % Partially Relevant 11.7 %
Groundedness 96.4 % Non-Relevant 16.6 %
Table 5. Expert evaluation of GPT-4o nugget quality and nugget-document relevance judgments on LangChain.
5.2.1. Nugget Quality Evaluation
For nugget quality evaluation, we ask the annotator to answer the following questions (A, B, and C) after reading the Stack Overflow question, answer, and list of nuggets. (1) $A$ does the nugget produce hallucinated content? requiring a boolean response (2) $B$ is the information provided in the nugget minor or redundant? also requiring a boolean response. After annotating all nuggets, we ask (3) $C$ how many additional nuggets are required to cover all key ideas, requiring an integer in the response.
Evaluation Metrics. After answering all the above questions, we measure quality by calculating three metrics: (1) precision: whether nuggets generated are accurate, (2) recall: whether nuggets cover the key aspects of the answer and (3) groundedness: whether nuggets produce non-hallucinated content, i.e., outside the scope of the answer. More formally, we define them as follows:
(2) $\text{Precision}=\dfrac{|Nuggets|-\mathrm{sum}(B)}{|Nuggets|}$
(3) $\text{Recall}=\dfrac{|Nuggets|-\mathrm{sum}(B)}{|Nuggets|-\mathrm{sum}(B)+C}$
(4) $\text{Groundedness}=\dfrac{|Nuggets|-\mathrm{sum}(A)}{|Nuggets|},$
where $|Nuggets|$ denotes the count of nuggets for a given question.
Experimental Results. As shown in Table 5, nuggets in the topic of LangChain achieve above 90% in precision and 96% in recall and groundedness, indicating GPT-4o can generate high-quality nuggets required in the FreshStack framework. Most nuggets are well-grounded, i.e., do not produce hallucinated content (3.6% error), and cover the key aspects of the answer in terms of recall (3.4% error). Precision errors are higher (9.9% error), showing nuggets may contain either minor or repeated information. Within these errors, the last positioned nugget is not informative in almost 50% of the time, and either the first or second positioned nugget in the rest of the error cases.
5.2.2. Relevance Judgment Quality Evaluation
We assess the relevance between nuggets and documents in nugget-level support. Since judging all documents (including negatives) for each nugget is cumbersome with a limited budget, we qualitatively evaluate the relevant pairs. We sample one positive document per question, totaling 60 randomly sampled nugget-document pairs. The annotator labels the relevance on a three-level scale: relevant, partially relevant, and non-relevant.
Experimental Results. As shown in Table 5, 71.7% of the judged nuggets and documents are relevant, including an additional 11.7% which are labeled partially relevant, indicating a high precision in GPT-4o support judgment. On the other hand, GPT-4o makes a mistake in judgment for 16.6% of the total questions. This discrepancy arises from several factors: some documents are relevant to only part of the nugget’s information, leading to mislabeling; ambiguity within the nugget content can cause misjudgments; and occasionally, literal grounding of a document in the nugget does not translate to semantic relevance in answering the question.
Figure 3. Main experiment results with retrieval and reranker baselines (including fusion) on five topics. Best scores in the Oracle setting are taken from Table 4 (Stack Overflow Nuggets in blue, Stack Overflow Answer in red). The reranker is the Voyage AI rerank-2 model (AI, 2024a). Plots show $\alpha$-nDCG-10, Coverage20, and Recall@50 scores for each topic (top to bottom).
6. Main Experiments
In this section, we evaluate leading neural retrievers and rerankers for the document retrieval setting on the FreshStack dataset, addressing RQ3 posed in our introduction. All models are evaluated out-of-the-box, i.e., using only the Stack Overflow question to retrieve documents, and do not include any information about the Stack Overflow answer or nuggets, ensuring a fair assessment.
Experimental Settings. We evaluate the retrieval models used as baselines during pooling in FreshStack: BM25, BGE (Gemma-2), E5 Mistral 7B, Voyage-large-2, and the ensemble fusion. In addition, we evaluate the Voyage AI rerank-2 (AI, 2024a) as the reranker with a 16K context length, reranking the top 50 documents retrieved from the first-stage retrieval system. Metrics used for evaluation are defined in Section 4.3: $\alpha$-nDCG@10, Coverage@20, and Recall@50.
6.1. Model Evaluation Results
Figure 3 shows the results for each retrieval baseline without and with VoyageAI rerank-2 reranking queries on each topic. Each model category is color-coded. We plot the $\alpha$-nDCG@10, Coverage@20, and Recall@50 scores from left to right. Key takeaways and findings are discussed below:
Accuracy gap between Oracle indicates plenty of headroom. Techniques from the Oracle setting (using Stack Overflow answers or nuggets) achieve a substantially higher $\alpha$-nDCG@10, Coverage@20, and Recall@50 in contrast to all models, including ensemble fusion and reranking with VoyageAI rerank-2. This highlights the complexity of answering FreshStack questions and demonstrates the headroom for improvement in existing code-based retrieval models to decrease the gap. Further research should explore techniques to improve models on specialized topics, containing challenging and long questions in the FreshStack dataset, instead of overfitting on existing academic IR benchmarks such as BEIR (Thakur et al., 2021).
Ensemble fusion outperforms individual models. Individual retrieval models demonstrate limited success on the FreshStack dataset; whereas, the ensemble fusion of four retrieval models outperforms each retrieval model across all metrics ($\alpha$-nDCG@10, Coverage@20, and Recall@50) and all five topics, except $\alpha$-nDCG@10 on Godot4. This highlights a crucial point in retrieval systems: a compound retrieval system (Zaharia et al., 2024), developed as an ensemble of retrieval models or something similar, is required to retrieve documents for niche and challenging topics, as an individual retriever is not sufficient, at present. While the ensemble fusion of retrieval systems achieves the best scores on the FreshStack dataset, their inefficiency at inference time adds up due to individual model inference, requiring alternatives.
Opportunities to improve reranking. When using weak first-stage retrieval, neural rerankers typically improve document ranking (Thakur et al., 2021), although it has been recently shown that this is not always the case when strong first-stage retrieval is used (Zamani et al., 2022; Jacob et al., 2024). Consistent with these recent observations, reranking provides benefits over BM25 for all topics in the FreshStack dataset. However, for our dense retrievers, reranking provides a clear benefit on some but not all datasets. Specifically, while the reranker enhances $\alpha$-nDCG@10, Coverage@20, and Recall@50 for LangChain, Yolo v7 & v8, and Godot4, it reduces those metrics on Laravel 10 & 11 and Angular 16, 17 & 18 for the dense retrievers. We suspect the reranker is better in certain programming languages such as Python, and we keep it as future work to understand the limitations of the neural reranker (Jacob et al., 2024). Separately, we also tried GPT-4o-mini (OpenAI, 2024) naively as a zero-shot listwise reranker following Sun et al. (2023) and found preliminary results underperformed heavily. We suspect this was due to formatting challenges when concatenating multiple long documents, and we leave the exploration of LLMs as rerankers (perhaps with fine-tuning) to future work.
7. Discussion
FreshStack is a general framework for building challenging IR datasets. We apply the framework to community-sourced questions (with curated answers) and documents sourced from GitHub repositories. The framework is adaptable to other domains like Stack Exchange or internal forums. While we benchmark a few key retrieval models, future work would continue to benchmark code-focused retrieval models like voyage-3 (AI, 2024b), Code-T5+ (Wang et al., 2023), CodeRankEmbed (Suresh et al., 2024), and Jina-Code-v2 (Günther et al., 2023), and rerankers such as CodeRankLLM (Suresh et al., 2024).
Answer Evaluation. We focused on the evaluation of the retrieval setting primarily due to two reasons: (1) Existing RAG datasets evaluate retrieval using relevance criteria only, however, we evaluated models based on both diversity and relevance criteria, and (2) a crucial step in FreshStack is sourcing and building a document corpus and developing a general framework for high-quality pools and automatic judgments, which we can evaluate better in the retrieval setting. As we have curated Stack Overflow answers available, future work can extend on evaluating the quality of LLM-based answer generation with metrics such as nugget-based recall, which calculates how many nuggets are supported within a system’s response. Given the high difficulty of retrieval, validated with results observed in Section 6, we anticipate answer generation by LLM to be similarly challenging.
Benchmark Contamination. The FreshStack dataset builds on Stack Overflow data and is susceptible to future data contamination. To address this, FreshStack is intentionally designed to be generic and extensible to new domains and languages. To mitigate data contamination, the framework can add newer questions in existing topics, retire old and contaminated topics, and add newer topics that develop in the future. The relevance of FreshStack in the community relies on a continued commitment to keeping it updated in the upcoming years.
8. Conclusion
The emergence of RAG has improved modern retrieval systems by allowing real-time data incorporation into LLMs. However, existing IR and RAG benchmarks that measure retrieval quality are outdated. In this work, we introduce a holistic framework, FreshStack, to construct challenging datasets to evaluate retrieval systems realistically. We source real user questions and answers from Stack Overflow and build a document corpus using technical documents from public GitHub repositories. Using FreshStack, we construct datasets on five niche topics and evaluate four frontier retrieval models and a reranker model in the document retrieval setting. The accuracy gap observed between the retrieval models and approaches in the Oracle setting indicates plenty of headroom for improvement, and we identify cases that may motivate future research in reranking. We hope FreshStack will encourage the community to build more challenging and realistic IR and RAG datasets in the future.
Acknowledgements.
We thank Sean Kulinski and Alexander Trott for helping us set up the grading notes prompt required in nugget generation. We also thank Jacob Portes, Max Marion, Matei Zaharia, and others from Databricks who provided feedback at the early stages of the project.